OpenStudy (anonymous):

What is the graph for these systems of equations? 4x + y = 5 3x – 7y = 27

OpenStudy (anonymous):

Set up as y=mx+b

OpenStudy (anonymous):

Solve each of the equations for y 4x + y = 5 (subtract 4x from each side) y = -4x + 5 So the slope is -4/1 and the y-intercept is 5 ----------- 3x – 7y = 27 (subtract 3x from each side) -7y = -3x + 27 (divide each side by -7) y = 3/7x + (-27/7) So the slope is 3/7 and the y- intercept is -27/7 So now when you graph it, you find the y point (5) and apply the slope to it.  do you understand so far?